OpenStudy (anonymous):

Please help (cos(x))/(1+sin(x))=1-sin(x) Solve the equation Multiplying both sides by 1+sin(x), using Pythagorean identities, and quadratic formula got me cos(x)=0 or cos(x)=1 Giving me the specific solutions, on an interval of [0,2pi], 0, pi/2,3pi/2, and 2pi Yet when I checked it graphically with the original equation 3pi/2 was not an answer Please help

OpenStudy (anonymous):

@jdoe0001 any thoughts?

OpenStudy (jdoe0001):

\(\bf \cfrac{cos(x)}{1+sin(x)}=1-sin(x)\ right?\) checking :)

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

And then multiplying by (1+sin(x)) on the RHS is difference of squares so that gets me 1-sin^2(x) which is then cos^2(x)=cos(x)

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

I think I see the issue your solution is correct however, when taking the inverse trig functions, recall their Range \(\textit{Inverse Trigonometric Identities} \\ \quad \\ \begin{array}{cccl} Function&{\color{brown}{ Domain}}&{\color{blue}{ Range}}\\ \hline\\ {\color{blue}{ y}}=sin^{-1}({\color{brown}{ \theta}})&-1\le {\color{brown}{ \theta}} \le 1&-\frac{\pi}{2}\le {\color{blue}{ y}}\le \frac{\pi}{2} \\ \quad \\ {\color{blue}{ y}}=cos^{-1}({\color{brown}{ \theta}})&-1\le {\color{brown}{ \theta}} \le 1& 0 \le {\color{blue}{ y}}\le \pi \\ \quad \\ {\color{blue}{ y}}=tan^{-1}({\color{brown}{ \theta}})&-\infty\le {\color{brown}{ \theta}} \le +\infty &-\frac{\pi}{2}\le {\color{blue}{ y}}\le \frac{\pi}{2} \end{array}\) notice, the Range for cosine

OpenStudy (jdoe0001):

they all three run on those constraints, thus

OpenStudy (anonymous):

Ohh I see Thanks!